If a random variable X follows Normal distribution such that P (9.65<=X<=13.8)=0.07008and P(X>9.6)=0.8159 where Find the mean & variance of X and find P(X>=13.8 /X>=9.6).


1
Expert's answer
2020-12-29T16:03:17-0500

P(X>9.6)=0.8159

P(x>z)=0.8159

z=-0.9


P(X>13.8) = P(X>9.6) - P (9.6"\leq" X"\leq" 13.8) = 0.8159 - 0.07008 = 0.7458

P(x>z)=0.7458

z=-0.661

z="\frac{X-\mu}{\sigma}"


-0.9="\frac{9.6-\mu}{\sigma}"

-0.661="\frac{13.8-\mu}{\sigma}"


"-0.9*\sigma+\mu=9.6"

"-0.661*\sigma+\mu=13.8"


"\sigma=17.6"

"\mu=25.4"


P(X"\ge"13.8|X"\ge"9.6) = P(X"\ge"13.8)/P(X"\ge"9.6) = 0.7458/0.8159 = 0.914